1.   A box at rest has the shape of a cube 2.6 m on a side. This box is loaded onto the flat floor of a spaceship and the spaceship then flies past us with a horizontal speed of 0.80c. What is the volume of the box as we observe it?



1
Expert's answer
2022-04-04T09:04:46-0400

Answer

the volume of the box as we will observe it as

"V=V_0\sqrt{1-\frac{v^2}{c^2}}\\=(2.6) ^3*\sqrt{1-\frac{(0.8c)^2}{c^2}}\\=2.6*2.6*2.6*0.06\\=1.05456m^3"